# M2CVD: Multi-Model Collaboration for Code Vulnerability Detection

### v1

@@ Abstract.
Large Language Models (LLMs) have strong capabilities in code comprehension, but fine-tuning costs and semantic alignment issues limit their project-specific optimization; conversely, code models such CodeBERT are easy to fine-tune, but it is often difficult to learn vulnerability semantics from complex code languages. To address these challenges, this paper introduces the Multi-Model Collaborative Vulnerability Detection approach (M2CVD) that leverages the strong capability of analyzing vulnerability semantics from LLMs to improve the detection accuracy of code models. M2CVD employs a novel collaborative process: first enhancing the quality of vulnerability semantic description produced by LLMs through the understanding of project code by code models, and then using these improved vulnerability semantic description to boost the detection accuracy of code models. We demonstrated M2CVD’s effectiveness on two real-world datasets, where M2CVD significantly outperformed the baseline. In addition, we demonstrate that the M2CVD collaborative method can extend to other different LLMs and code models to improve their accuracy in vulnerability detection tasks.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Vulnerabilities in software refer to code weaknesses that can be easily exploited, which can lead to serious consequences such as unauthorized information disclosure  (Fu and Tantithamthavorn, 2022) and cyber extortion  (Thapa et al., 2022). Recent statistics underscore this burgeoning issue: In Q1 of 2022, the US National Vulnerability Database (NVD) disclosed 8,051 vulnerabilities, marking a 25% surge from the previous year  (Cheng et al., 2022). Further accentuating this trend, a study revealed that out of 2,409 analyzed codebases, 81% had at least one recognized open-source vulnerability. The increasing scale and ubiquity of these vulnerabilities emphasize the need for well-developed automated vulnerability detection mechanisms. Such a detection system helps to strengthen software security and forestall a range of potential security risks (Fu and Tantithamthavorn, 2022; Thapa et al., 2022; Jang-Jaccard and Nepal, 2014; Johnson et al., 2011).
The vulnerability detection models in the existing literature are mainly divided into two categories: (1) conventional detection models  (Yamaguchi, 2015, 2017) and (2) Deep Learning(DL)-based models  (Li et al., 2021a; Duan et al., 2019; Russell et al., 2018; Li et al., 2018; Dam et al., 2017) The former typically requires experts to manually formulate detection rules (Checkmarx, 2022; Flawfinder, 2022). These methods are usually labor-intensive to create and are difficult to achieve low false positive rates and low false negative rates  (Li et al., 2018, 2021b). On the contrary, deep learning (DL) -based detection methods learn the patterns of vulnerabilities from a training set (Li et al., 2018; Dam et al., 2017; Lin et al., 2017). They avoid manual heuristic methods and autonomously learn and identify vulnerability features. In order to further learn the semantics of vulnerabilities, methods based on pre-trained code models (Nguyen et al., 2022) and vulnerability detection studies based on LLMs (Fu et al., 2023) have been proposed successively. In summary, traditional vulnerability detection methods are usually depend on pre-defined rules, a process of expert interventions, rendering them laborious and occasionally imprecise. The DL-based detection methods, by contrast, show better detection ability by learning vulnerable code patterns automatically.
In the latest research (Steenhoek et al., 2023a), the efficacy of pre-trained language models for software vulnerability detection has been extensively explored, encompassing LLMs such as ChatGPT (openAI, 2022) and LLaMa (Llama, 2022), alongside pre-trained code models like CodeBERT (Feng et al., 2020) and UniXcoder (Guo et al., 2022). Compared with traditional deep learning networks, these models show more excellent performance in code vulnerability detection tasks after fine-tuning (Nguyen et al., 2022).
@@
In contrast to existing approaches that use pattern matching to enhance vulnerability semantic  (Steenhoek et al., 2023b), we resort tothe strong understanding abilities of LLMs to create natural language descriptions of vulnerable code, so as to make connections between code and the causes of vulnerabilities. This will bring two benefits: the more abstract natural language description will help the code model better learn the semantics of the vulnerability, and the vulnerability description can help programmers better determine the cause of the vulnerability to maintain the code. But we’ve hit the second challenge, the semantic alignment problem of LLMs. Given the scale of LLMs, fine-tuning them on a specific domain is tough. Real projects, organizations, or specific fields have their own coding rules and business logic. Using LLMs trained on data from open domain might not make accurate vulnerability judgment on code in a specific domain (e.g., Fu et al. (2023) reports a F1 score of 29% with GPT4). As a result, LLMs may generate incorrect vulnerability description.
In this paper, we’ve proposed M2CVD, an innovative approach that combines the strengths of pre-trained code models and LLMs to better detect vulnerabilities. For the first challenge, we rely on the ability of LLMs to interpret vulnerabilities, leveraging the explanatory text to hellp code models to understand the semantics of vulnerabilities. For the second challenge, we rely on the advantage of code models that is easy to fine-tune, and use their judgment results to enhance the vulnerability semantic understanding of LLMs for specific projects. In this way, M2CVD can help operators to improve the accuracy of vulnerability detection through the collaborative interaction process combined with LLMs API without changing the existing code model structure. In summary, the main contributions of this paper are as follows:
@@
    We evaluate our approach through extensive experimentation on two real-world datasets. Experimental results show that the M2CVD can still improve the performance of code vulnerability detection with the different of pre-trained code model and LLMs.  [^1] Our anonymous replication package (data and code) :https://figshare.com/articles/dataset/M2CVD/24831567.
@@ 3. Approach
In general, M2CVD requires three models to work together, including code model $f_{d}$,$f_{v}$ and LLM $f_{c}$, and relies on the collaborative interaction of $f_{d}$ and $f_{c}$ to assist the enhancement of $f_{v}$ model in the vulnerability detection task. The overall framework of M2CVD as shown in 1, consisting of three phases:
@@ 3.3. Integrated Vulnerability Detection
In Phase III, M2CVD leverages the vulnerability assessments and vulnerability descriptions from the LLM to supplement the vulnerability semantics of the vulnerability code. Initially, we restructure the input dataset $L$. such that $L_{c}=(P,C,N)$. The $C$ represents ChatGPT’s assessment, while $N$ denotes the LLM’s description of the vulnerability in the code segment. After integrating the vulnerability semantics into the dataset $L_{c}=(P,C,N)$, the validation model $f_{v}$ was obtained by fine-tuning.
@@
For the fine-tuning process of the validation model, it is crucial to ensure that the training set undergoes both Phase I and II for filling in vulnerability semantics. This approach differs significantly from merely using a vulnerability label for semantic completion via the LLM. We have found that limiting the enhancement of code vulnerability semantics to only those with identified vulnerabilities in the training set can lead to overfitting in the validation model. Consequently, to adequately supplement the entire training set’s semantic content, historical data plays a pivotal role in the vulnerability semantic generation process during Phases I and II. This necessity occasionally leads ChatGPT to erroneously add semantics to codes perceived as vulnerable. Nevertheless, maintaining the same data source during both the training and inference phases, while integrating a degree of noise, has effectively increased the robustness of the validation model.
@@ 4. Experiments
Table 1. Comparison results for different models on Devign and Reveal datasets.The best result for each metric is highlighted in bold. The cells in grey represent the performance of the top-3 best methods in each metric, with darker colors representing better performance.
Dataset Devign (Zhou et al., 2019) Reveal (Chakraborty et al., 2021)
Models Accuracy Recall Precision F1 Accuracy Recall Precision F1
ChatGPT3.5-COT 49.83 32.24 33..00 30.61 63.72 26.34 30.54 27.70
Devign 56.89 52.50 64.67 57.59 87.49 31.55 36.65 33.91
SySeVR 47.85 46.06 58.81 51.66 74.33 40.07 24.94 30.74
ReGVD 61.89 48.20 60.74 53.75 90.63 14.47 64.70 23.65
CodeBERT 63.59 41.99 66.37 51.43 91.33 34.64 62.20 44.50
UniXcoder 64.82 54.48 63.53 58.94 91.90 26.31 78.94 39.47
CodeT5 65.04 54.26 64.12 58.78 88.79 38.16 43.28 40.56
M2CVD 68.11 54.50 69.51 61.09 92.12 26.75 83.56 40.56
@@ 4.1. Datesets
To evaluate the effectiveness of M2CVD, we employ two datasets from real projects :(1) Devign  (Zhou et al., 2019), and (2) Reveal (Chakraborty et al., 2021). The Devign dataset, derived from a graph-based code vulnerability detection study  (Zhou et al., 2019), stands as a dataset of function-level C/C++ source code from the well-established open-source projects QEMU and FFmpeg. Aligning with the methodology articulated by Li et al.  (Zhou et al., 2019), the partitioning of the Devign dataset adheres to a conventional 80:10:10 ratio, demarcating the bounds for training, validation, and testing data, respectively. The dataset completes the labeling of vulnerable code by a group of security researchers performing a rigorous two-stage review. In the task of software vulnerability detection, the REVEAL dataset is a representative dataset, as presented in  (Chakraborty et al., 2021). It is a further exploration of data redundancy and unrepresentative class distributions in existing datasets. As a detection code dataset, REVEAL encompasses source code extracted from two open-source forays: the Linux Debian kernel and Chromium. Similar to the real-world situation, this dataset has an imbalanced label distribution, with the number of normal code fragments much larger than the number of vulnerable ones (10:1). Similarly, in the Reveal dataset, a split ratio of 80:10:10 was set.
@@ 4.2. Performance Metrics
In the process of evaluating the performance of the model, the proposed method includes four metrics widely recognized in the field of software testing and analysis(Zhou et al., 2019):
@@ 4.3. Baseline Methods
(4) SySeVR (Li et al., 2021b): SySeVR uses code statements, program dependencies, and program slicing as capabilities, and utilizes bidirectional recurrent neural networks to detect vulnerable code fragments.
@@
(7) UniXcoder (Guo et al., 2022): This is a unified code representation model that leverages a Transformer-based architecture. UniXcoder extends the capabilities of models like CodeBERT by incorporating a comprehensive understanding of code syntax and semantics, thereby enhancing the model’s performance in coding tasks such as code summarization, translation, and completion.
Experiment Environment: We implemented M2CVD in Python using Tensorflow. The experiments were performed on a machine containing a NVIDIA GeForce GTX A6000 Gpu and two Intel Xeon Gold 6226R 2.90 GHz cpus.
@@ 5. EXPERIMENTS
In this section, we conduct extensive experiments to justify our model’s superiority and analyze the reasons for its effectiveness. Specifically, we aim to answer the following research questions:
@@
RQ3: What are the effects of vulnerability description refinement for LLMs?
RQ4: What are the effects of different pre-trained code models and LLMs working together on vulnerability detection?
@@
To answer the first question, we compare M2CVD with the seven baseline methods on the two datasets as shown in table 2. We can draw conclusions about the performance of M2CVD compared to the baselines across the evaluated datasets. Here is a comprehensive performance analysis for M2CVD:
@@
M2CVD demonstrates a marked superiority in terms of Accuracy on both datasets. In the Devign dataset, M2CVD attains the highest Accuracy of 68.11%, the highest F1 score of 61.09% and the highest Precision outperforming all other models. This indicates that M2CVD has the most balanced performance in correctly identifying vulnerabilities without being skewed towards over-predicting (which would increase recall but decrease precision) or under-predicting (which would do the opposite).
@@
On the Reveal dataset, M2CVD again stands out with the highest Accuracy of 92.12%. Both Precision and F1 metrics maintain the level of optimal level. These figures not only show that M2CVD maintains its high performance in different testing conditions but also that it consistently understands and predicts code vulnerabilities with high precision and recall. The performance improvement of M2CVD on the reveal dataset is much less than that of Devign, which we believe is caused by the imbalance of the Reveal dataset, and the vulnerability data is far less than the normal data. This allows ChatGPT to add far less vulnerability semantics to this dataset than to the Devign dataset. Experimental reality, ChatGPT has ultra-high recall on both datasets. After case analysis, we found that ChatGPT tends to think that the code does not have vulnerabilities when the code context is not provided.
@@
It is worth noting that the underlying code model for M2CVD is the UniXcoder. After the multi-model collaboration process, M2CVD demonstrates superior accuracy and precision over UniXcoder, with marked improvements seen in both the Devign and Reveal datasets. Overall, the performance of M2CVD shows that M2CVD has a more balanced and higher performance in overall performance compared to UniXcoder.
@@
Result Answer to RQ1: The performance of M2CVD shows that compared with a single model, M2CVD effectively achieves higher performance in code defect detection tasks under different experimental conditions through a collaborative mechanism.
@@
M2CVD(CodeBERT) w/o PII 65.50%
M2CVD(CodeBERT) 66.10%
UniXcoder 64.82%
M2CVD(UniXcoder) w/o PII 67.05%
M2CVD(UniXcoder) 68.11%

### v2

# M2CVD: Enhancing Vulnerability Semantic through Multi-Model Collaboration for Code Vulnerability Detection

@@ Abstract.
Large Language Models (LLMs) have strong capabilities in code comprehension, but fine-tuning costs and semantic alignment issues limit their project-specific optimization; conversely, code models such as CodeBERT are easy to fine-tune, but it is often difficult to learn vulnerability semantics from complex code languages. To address these challenges, this paper introduces the Multi-Model Collaborative Vulnerability Detection approach (M2CVD) that leverages the strong capability of analyzing vulnerability semantics from LLMs to improve the detection accuracy of code models. M2CVD employs a novel collaborative process: first enhancing the quality of vulnerability semantic description produced by LLMs through the understanding of project code by code models, and then using these improved vulnerability semantic descriptions to boost the detection accuracy of code models. M2CVD include three main phases: 1) Initial Vulnerability Detection: The initial vulnerability detection is conducted by fine-tuning a detection model (e.g., CodeBERT) and interacting with a LLM (e.g., ChatGPT) respectively. The vulnerability description will be generated by the LLM when the code is detected vulnerable by the LLM. 2) Vulnerability Description Refinement: By informing the LLM of the vulnerability assessment results of the detection model, we refine the vulnerability description by interacting with the LLM. Such refinement can enhance LLM’s vulnerability understanding in specific projects, effectively bridging the previously mentioned alignment gap; 3) Integrated Vulnerability Detection: M2CVD integrates code fragment and the refined vulnerability descriptions inferred to form synthetic data. Then, the synthetic data is used to fine-tune a validation model, optimize the defect feature learning efficiency of the model, and improve the detection accuracy. We demonstrated M2CVD’s effectiveness on two real-world datasets, where M2CVD significantly outperformed the baseline. In addition, we demonstrate that the M2CVD collaborative method can extend to other different LLMs and code models to improve their accuracy in vulnerability detection tasks.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Vulnerabilities in software refer to code weaknesses that can be easily exploited, which can lead to serious consequences such as unauthorized information disclosure  (Fu and Tantithamthavorn, 2022) and cyber extortion  (Thapa et al., 2022). Recent statistics underscore this burgeoning issue: In Q1 of 2022, the US National Vulnerability Database (NVD) disclosed 8,051 vulnerabilities, marking a 25% increase from the previous year  (Cheng et al., 2022). Further accentuating this trend, a study revealed that out of 2,409 analyzed codebases, 81% had at least one recognized open-source vulnerability. The increasing scale and ubiquity of these vulnerabilities emphasize the need for well-developed automated vulnerability detection mechanisms. Such a detection system helps to strengthen software security and forestall a range of potential security risks (Fu and Tantithamthavorn, 2022; Thapa et al., 2022; Jang-Jaccard and Nepal, 2014; Johnson et al., 2011).
The vulnerability detection models in the existing literature are mainly divided into two categories: (1) conventional detection models  (Yamaguchi, 2015, 2017) and (2) Deep Learning (DL)-based models  (Li et al., 2021a; Duan et al., 2019; Russell et al., 2018; Li et al., 2018; Dam et al., 2017) The former typically requires experts to manually formulate detection rules (Checkmarx, 2022; Flawfinder, 2022). These methods are usually labor-intensive to create and are difficult to achieve low false positive rates and low false negative rates  (Li et al., 2018, 2021b). On the contrary, deep learning (DL) -based detection methods learn the patterns of vulnerabilities from a training set (Li et al., 2018; Dam et al., 2017; Lin et al., 2017). They avoid manual heuristic methods and autonomously learn and identify vulnerability features. In order to further learn the semantics of vulnerabilities, methods based on pre-trained code models (Nguyen et al., 2022) and vulnerability detection studies based on LLMs (Fu et al., 2023) have been proposed successively. In summary, traditional vulnerability detection methods usually depend on pre-defined rules, a process of expert interventions, rendering them laborious and occasionally imprecise. The DL-based detection methods, by contrast, show better detection ability by learning vulnerable code patterns automatically.
In the latest research (Steenhoek et al., 2023a), the efficacy of pre-trained language models for software vulnerability detection has been extensively explored, encompassing LLMs such as ChatGPT (openAI, 2022) and LLaMa (Llama, 2022), alongside pre-trained code models like CodeBERT (Feng et al., 2020) and UniXcoder (Guo et al., 2022). Compared with traditional deep learning networks, these models show more excellent performance in code vulnerability detection tasks after fine-tuning (Nguyen et al., 2022). Figure 1 illustrates the flow of code model fine-tuning.
@@
In contrast to existing approaches that use pattern matching to enhance vulnerability semantic  (Steenhoek et al., 2023b), we resort to the strong understanding abilities of LLMs to create natural language descriptions of vulnerable code, so as to make connections between code and the causes of vulnerabilities. This will bring two benefits: the more abstract natural language description will help the code model better learn the semantics of the vulnerability, and the vulnerability description can help programmers better determine the cause of the vulnerability to maintain the code. In the latest research, this process is exploited for natural language understanding(Min et al., 2024). But we’ve hit the second challenge, the semantic alignment problem of LLMs. Given the scale of LLMs, fine-tuning them on a specific domain is challenging. Real projects, organizations, or specific fields have their own coding rules and business logic. Using LLMs trained on data from open domain might not make accurate vulnerability judgment on code in a specific domain (e.g., Fu et al. (2023) reports a F1 score of 29% with GPT4). As a result, LLMs may generate incorrect vulnerability description.
Figure 1. Existing vulnerability detection method processes based on code model fine-tuning and interactive vulnerability detection processes based on large language models.
In this paper, we have proposed M2CVD, an innovative approach that combines the strengths of pre-trained code models and LLMs to better detect vulnerabilities. For the first challenge, we rely on the ability of LLMs to interpret vulnerabilities, leveraging the explanatory text to help code models to understand the semantics of vulnerabilities. For the second challenge, we rely on the advantage of code models that is easy to fine-tune, and use their judgment results to enhance the vulnerability semantic understanding of LLMs for specific projects. In this way, M2CVD can help operators to improve the accuracy of vulnerability detection through the collaborative interaction process combined with LLMs API without changing the existing code model structure. In summary, the main contributions of this paper are as follows:
@@
    We evaluate our approach through extensive experimentation on two real-world datasets. Experimental results show that the M2CVD can still improve the performance of code vulnerability detection with the different of pre-trained code model and LLMs.
Data Availability. We open-source our replication package [^1] Our replication package (data and code) :https://github.com/HotFrom/M2CVD, including the datasets and the source code of M2CVD, to facilitate other researchers and practitioners to repeat our work and verify their studies.
Paper Organization. Section 2 describes the background of code vulnerability detection. Section 3 presents our model M2CVD. Sections 4 and 5 describe the datasets and experiments of our study, respectively. Sections 6 discuss a case of M2CVD, respectively. Section 7 concludes the discussion of M2CVD. And Section 8 includes summary of our approach and future directions.
@@ 3. Approach
In general, M2CVD requires three models to work together, including code model $f_{d}$,$f_{v}$ and LLM $f_{c}$, and relies on the collaborative interaction of $f_{d}$ and $f_{c}$ to assist the enhancement of $f_{v}$ model in the vulnerability detection task. The overall framework of M2CVD as shown in Figure 2, consisting of three phases:
@@
In the default configuration of this paper, the code model used by M2CVD is UnixCoder and LLM is ChatGPT 3.5.
@@ 3.3. Integrated Vulnerability Detection
In Phase III, M2CVD leverages the vulnerability assessments and vulnerability descriptions from the LLM to supplement the vulnerability semantics of the vulnerability code. Initially, we restructure the input dataset $L$. such that $L_{c}=(P,C,N,Y)$. The $C$ represents LLM’s assessment, while $N$ denotes the LLM’s description of the vulnerability in the code segment. After integrating the vulnerability semantics into the dataset $L_{c}=(P,C,N,Y)$, the validation model $f_{v}$ was obtained by fine-tuning.
@@
Data Identically Distributed Guarantee. For the fine-tuning process of the validation model, it is crucial to ensure that the training set undergoes both Phase I and II for filling in vulnerability semantics. This approach differs significantly from merely using a vulnerability label for semantic completion via the LLM. We have found that limiting the enhancement of code vulnerability semantics to only those with identified vulnerabilities in the training set can lead to overfitting in the validation model. Therefore, in order to ensure that the training set and the validation test data remain identically distributed as much as possible, it is necessary to prohibit directly informing LLM data labels during the vulnerability semantic generation process in Phases I and II. This necessity occasionally leads ChatGPT to erroneously add semantics to codes perceived as vulnerable. Nevertheless, maintaining the same data source during both the training and inference phases, while integrating a degree of noise, has effectively increased the robustness of the validation model.
@@ 3.5. Loss Function
3.6. Implementation Details
M2CVD has two processes of model fine-tuning, in which the fine-tuning process of the first evaluation model adopts the best performance parameters reported by the existing code model, specifically referring to [^2] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/Defect-detection. After completing the vulnerability assessment report, M2CVD implements the second model fine-tuning process, in which epoch is 4, the sequence length takes the maximum length of the base code model used, which is 1024 for uniXcoder by default in this paper, the learning rate is 2e-5, and the bachsize is 12.
@@ 4.1. Datesets
To evaluate the effectiveness of M2CVD, we employ two datasets from real projects :(1) Devign  (Zhou et al., 2019), and (2) Reveal (Chakraborty et al., 2021). The Devign dataset, derived from a graph-based code vulnerability detection study  (Zhou et al., 2019), stands as a dataset of function-level C/C++ source code from the well-established open-source projects QEMU and FFmpeg. Aligning with the methodology articulated by Li et al.  (Zhou et al., 2019), the partitioning of the Devign dataset adheres to a conventional 80:10:10 ratio, demarcating the bounds for training, validation, and testing data, respectively. The dataset completes the labeling of vulnerable code by a group of security researchers performing a rigorous two-stage review. In the task of software vulnerability detection, the REVEAL dataset is a representative dataset, as presented in  (Chakraborty et al., 2021). It is a further exploration of data redundancy and unrepresentative class distributions in existing datasets. As a detection code dataset, REVEAL encompasses source code extracted from two open-source forays: the Linux Debian kernel and Chromium. Similar to the real-world situation, this dataset has an imbalanced label distribution, with the number of normal code fragments much larger than the number of vulnerable ones (10:1). Similarly, in the Reveal dataset, a split ratio of 80:10:10 was set (Ding et al., [n. d.]).
@@ 4.2. Performance Metrics
In the process of evaluating the performance of the model, the proposed method includes three metrics widely recognized in the field of software testing and analysis(Zhou et al., 2019):
@@
REVEAL is an imbalanced dataset, so we emphasize the use of F1 as the evaluation metric (Ding et al., [n. d.]). The designed dataset is balanced, so we follow the original benchmark to report the classification accuracy (Zhou et al., 2019).
Since the model performance can vary with different random seeds (Steenhoek et al., 2023a), we used the random seed setting commonly used in existing open source methods, seed=42 (Feng et al., 2020; Ding et al., [n. d.]).
@@ 4.3. Baseline Methods
(6) UniXcoder-base (Guo et al., 2022): This is a unified code representation model that leverages a Transformer-based architecture. UniXcoder extends the capabilities of models like CodeBERT by incorporating a comprehensive understanding of code syntax and semantics, thereby enhancing the model’s performance in coding tasks such as code summarization, translation, and completion.
(7) UniXcoder-base-nine: Continue pre-training unixcoder-base on NL-PL pairs of CodeSearchNet dataset and additional 1.5M NL-PL pairs of C, C++ and C# programming language. The model can support nine languages: java, ruby, python, php, javascript, go, c, c++ and c#.
(8) TRACED (Ding et al., [n. d.]): TRACED employs an execution-aware pre-training strategy to enhance code models’ understanding of dynamic code properties, significantly improving their performance in execution path prediction, runtime variable value prediction, clone retrieval, and vulnerability detection.
Experiment Environment: We implemented M2CVD in Python using Tensorflow. The experiments were performed on a machine containing three NVIDIA GeForce GTX A6000 GPU and two Intel Xeon Gold 6226R 2.90 GHz cpus.
@@ 5. Experiments
In this section, we conduct extensive experiments to demonstrate our model’s superiority and analyze the reasons for its effectiveness. Specifically, we aim to answer the following research questions:
@@
In this RQ, the performance of M2CVD is verified with two real-world vulnerability datasets. Given a code fragment to be detected, M2CVD generates a vulnerability description through a large model, and generates code and vulnerability assessment pairs through collaborative process changes. The performance of M2CVD on the test data is evaluated and compared to the SOTA baseline on two datasets.
@@
In this RQ, we verified the effectiveness of the components, which included the comparison between the results of fine-tuning after generating the vulnerability description directly through the large model and the results of fine-tuning after refining the vulnerability description in step 2.
RQ3: What are the effects of hints of code models for LLMs?
In this RQ, we verify whether the first judgment results of the code model have a positive effect on the LLM. We inform the LLM code model or error judgment information respectively to verify the rationality of the M2CVD process.
RQ4: How well does M2CVD generalize across different code models and LLMs?
In this RQ, we validate the performance between different code models and LLMs combinations on the vulnerability detection task, validating the generality of the M2CVD approach.
Table 1. Comparison results for different models on Devign and Reveal datasets. The best result for each metric is highlighted in bold.
Dataset Devign (Zhou et al., 2019) Reveal (Chakraborty et al., 2021)
Models Acc Recall Prec F1 F1 Recall Prec Acc
ChatGPT 3.5 COT 49.83 32.24 33.00 30.61 27.70 26.34 30.54 63.72
ChatGPT 4o COT 53.73 7.46 45.94 4.06 12.51 22.17 97.33 20.97
Devign 56.89 52.50 64.67 57.59 33.91 31.55 36.65 87.49
ReGVD 61.89 48.20 60.74 53.75 23.65 14.47 64.70 90.63
CodeBERT 63.59 41.99 66.37 51.43 35.11 25.87 54.62 90.41
UniXcoder-base 65.77 51.55 66.42 58.05 39.47 26.31 78.94 91.90
CodeT5-base 65.04 54.26 64.12 58.78 40.56 38.16 43.28 88.79
UniXcoder-nine 66.98 56.33 66.63 61.05 42.19 33.77 56.20 90.72
TRACED 64.42 61.27 60.03 61.05 32.66 21.49 68.05 91.11
M2CVD 68.33 57.76 68.39 62.63 48.10 39.03 62.67 91.55
*The Reveal dataset is an imbalanced dataset. F1 and recall are the primary metrics.
@@
To answer the first question, we compare M2CVD with the seven baseline methods on the two datasets as shown in table 1. We can draw conclusions about the performance of M2CVD compared to the baselines across the evaluated datasets.
Table 1 presents the performance of ChatGPT on the vulnerability detection task. It is evident that large-scale language models employ aggressive detection logic. Specifically, in ChatGPT 4o, nearly all code snippets were identified as vulnerable, leading to considerably low F1 scores across both datasets.
@@
M2CVD demonstrates a marked superiority in terms of Accuracy on both datasets. In the Devign dataset, M2CVD attains the highest Accuracy of 68.33%, the highest F1 score of 62.63% and the highest Precision outperforming all other models. This indicates that M2CVD has the most balanced performance in correctly identifying vulnerabilities without being skewed towards over-predicting (which would increase recall but decrease precision) or under-predicting (which would do the opposite).
@@
On the Reveal dataset, since the proportion of negative samples in this dataset is 90, the model with strong fitting performance generally exceeds 90% on ACC. In this dataset, people are generally interested in the ability of the model to find positive samples (vulnerability). For M2CVD, Both Recall and F1 metrics maintain the level of optimal level. These figures not only show that M2CVD maintains its high performance in different testing conditions but also that it consistently understands and predicts code vulnerabilities with high precision and recall. The performance improvement of M2CVD on the reveal dataset is much less than that of Devign, which we believe is caused by the imbalance of the Reveal dataset, and the vulnerability data is far less than the normal data. This allows ChatGPT to add far less vulnerability semantics to this dataset than to the Devign dataset.
Figure 3 presents a comparison of three models (M2CVD, TRACED, and UniXcoder) through Venn diagrams, highlighting their performance in detecting vulnerabilities and false negatives. In Figure 3(a), the Venn diagram illustrates the overlap in vulnerabilities correctly detected by the three models: M2CVD independently detects 92 vulnerabilities, outperforming TRACED, which detects 27, and UniXcoder, which detects 53. The three models collectively identify 856 vulnerabilities, indicating a significant overlap and suggesting that they are effective in detecting similar vulnerabilities. M2CVD and UniXcoder jointly detect 142 vulnerabilities, while M2CVD and TRACED jointly detect 72. Figure 3(b) depicts the distribution of false negatives among the three models: M2CVD independently produces 18 false negatives, which is significantly lower than the 142 produced by TRACED and 72 by UniXcoder. The three models collectively produce 217 false negatives, indicating some common challenges in detecting certain vulnerabilities. In conclusion, the analysis of Figure 3 demonstrates that the M2CVD model not only excels in detecting a higher number of actual vulnerabilities but also has a lower false negative rate compared to TRACED and UniXcoder. This indicates that the M2CVD model offers superior overall performance in vulnerability detection tasks, making it a valuable tool for applications requiring high accuracy and low false negative rates. Figure 6 shows the situation in the Reveal dataset. As shown in Figure Figure 6(a), M2CVD finds more vulnerabilities. However, M2CVD also misses 6 vulnerabilities than UniXcoder due to the complexity of imbalanced data.
@@
In order to better compare the performance of M2CVD, we adopt the UniXcoder-base as the base model, which has the same number of parameters as baseline models such as TRACED and CodeBERT. After the multi-model collaboration process, M2CVD demonstrates superior accuracy and precision over UniXcoder-base, with marked improvements seen in both the Devign and Reveal datasets. Overall, the performance of M2CVD shows that M2CVD has a more balanced and higher performance in overall performance compared to UniXcoder.
(a) Correct Predictions (Target = 0)
(b) False Negatives (Predicted = 0, True = 1)
Figure 3. Comparison of the performance of different models in dataset Devign. (a) shows the number of vulnerabilities detected by the M2CVD, TRACED, and UniXcoder models. (b) shows the number of missed vulnerabilities by each model.
@@
Answer to RQ1: The performance of M2CVD shows that compared with a single model, M2CVD effectively achieves higher performance in code defect detection tasks under different experimental conditions through a collaborative mechanism.
@@
M2CVD(GPT 3.5+CodeBERT) w/o PII 65.50%
M2CVD(GPT 3.5+CodeBERT) 66.10%
UniXcoder-base 64.82%
M2CVD(GPT 3.5+UniXcoder-base) w/o PII 67.05%
M2CVD(GPT 3.5+UniXcoder-base) 68.33%
M2CVD(GPT 4o+UniXcoder-nine) w/o PII 64.45%
M2CVD(GPT 4o+UniXcoder-nine) 69.11%
Figure 4. Effects of vulnerability description refinement for M2CVD. The figure shows the improvement of the detection effect after CodeBert and UnixCoder pass the two stages of M2CVD, respectively.
Figure 5. Effects of different code models and LLMs. The figure shows the detection performance improvement of CodeBert and UnixCoder by the vulnerability description models provided by different LLMS, respectively.
(a) Correct Predictions (Target = 0)
(b) False Negatives (Predicted = 0, True = 1)
Figure 6. Comparison of Model Predictions Result in Reveal dataset